OpenStudy (anonymous):

Solve: (x - 8)(x + 11) = 0 Solve the equation (4x + 1)(3x - 2) = 0

OpenStudy (anonymous):

1. (x - 8)(x + 11) = 0 => x-8=0 or x+11=0 x-8=0 x=8 x+11=0 x=-11 2. (4x + 1)(3x - 2) = 0 => 4x+1=0 or 3x-2=0 4x+1=0 x=-1/4 3x-2=0 x=2/3